Weston Corporation just paid a dividend of $1.25 a share (i.e., D0 = $1.25). The dividend is expected to grow 11% a year for the next 3 years and then at 4% a year thereafter. What is the expected dividend per share for each of the next 5 years? Do not round intermediate calculations. Round your answers to the nearest cent.

D1 = $  
D2 = $  
D3 = $  
D4 = $  
D5 = $  

PLEASE SHOW WORK PLEASE. SHOW HOW IN CALCULATOR OR BASIC MATH STRUCTURE IF POSSIBLE. THANK YOU
Answer #1
D1 (1.25*1.11)=$1.39(Approx).
D2 (1.39*1.11)=$1.54(Approx).
D3 (1.54*1.11)=$1.71(Approx).
D4 (1.71*1.04)=$1.78(Approx).
D5 (1.78*1.04)=$1.85(Approx)